# Reflectionless modes as a source of Weyl nodes in multiterminal Josephson junctions

### v1

@@ Abstract
Multiterminal Josephson junctions are a promising platform to study non-trivial topology in engineered quantum systems. Yet, experimentally meaningful insight into what exactly makes these systems topologically non-trivial remains elusive. In this work, we show that zero energy reflectionless scattering modes (RSMs) of the normal scattering matrix result in topological phase boundaries. By analyzing two different setups, we explain the origin of each topological phase boundary and furthermore provide generalizable insight into these systems. The considerations here can be of help for experimentalists as it connects the properties of the normal scattering region to the Andreev bound state spectrum of superconducting junctions in a multiterminal setup.
@@
Introduction.– Topology manifests itself as one of the most fundamental aspects of condensed matter systems. Such systems include topological insulators [1, 2, 3, 4, 5], topological superconductors [6, 7, 8, 9] and quantum Hall systems [10, 11, 12]. Another promising direction lies in the ability to engineer the topological properties via control parameters to study topolgical phase transistions. Examples include topological photonics [13, 14, 15], topological-driven Floquet systems [16, 17], topological electrical [18] and superconducting circuits [19, 20, 21, 22]. Multiterminal Josephson junctions (MTJJs) have proven to be another promising route towards engineered topology where the topology is encoded in the superconducting phase degree of freedom [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Additionally, MTJJs also host novel (topological) phases in the context of non-hermiticity [43, 44, 45, 46, 47, 48, 49, 50, 51]. Whereas observable consequences of the non-trivial topology are predicted to exist [23, 31], whether these signatures can be observed in real systems is not clear [52]. Experimentally, recent advances in sample fabrication lead to the spectroscopic measurement of Andreev bound states (ABSs) with phase control via flux lines [53] which also represents an experimental realization of an Andreev molecule [54, 55]. In particular, phase manipulation of a four-terminal MTJJ was established experimentally in Ref. [56], which satisfies the minimal requirement for the existence of Weyl nodes in MTJJs with time-reversal symmetry [23]. Whereas the theory provided in Ref. [56] predicts non-trivial topology in accordance with spectroscopic data, additional experimental evidence is needed to proof the existence of non-trivial topology.
@@
Although MTJJs and their topology have been studied in detail for nearly a decade, making a connection from non-trivial topology to experimentally relevant parameters is difficult. Generally, promising samples must host ABSs very close to zero energy, yet in a multiterminal setup, such a zero energy ABS is not necessarily related to a unity transmission eigenmode as it is not possible to label ABSs by transmission eigenvalues of the full scattering matrix [57].
@@
In this manuscript, we aim to provide a concrete explanation for the appearance of non-trivial topology and Weyl nodes in MTJJs by demonstrating that reflectionless scattering modes (RSMs) [58] of the normal scattering matrix result in topological phase boundaries. We illustrate our findings by firstly analyzing a four-terminal hybrid system used in Ref. [59] where we discuss the conditions for topological phase boundaries in detail. We show that each phase boundary is related to a zero-energy RSM (zero-RSM) of the normal scattering matrix. In the second part, we analyze the four-terminal system used in Ref. [56] and its corresponding topological phase diagram in detail. We find that most topological phase boundaries correspond to zero-RSMs of the normal scattering matrix. Additionally, we find abnormal topological phase boundaries, which do not arise from RSMs, yet are related to effective reflectionless modes of a lower dimensional system. Additionally, we find that the source of topology is inherent to the non-trivial topology of RSMs [58]. The findings here can help provide explanations for the occurrence of zero-energy solutions and potentially topology in experimental MTJJ setups.
Reflectionless scattering modes (RSMs).– We consider a normal scattering region parametrized by a $n\times n$ normal scattering matrix $S_{\rm N}$. The normal region is coupled to $n$ superconducting (SC) leads with respective phases $\phi_{i}$, $\bm{\phi}={\rm diag}(\phi_{1},...,\phi_{n})$. Let’s assume that $S_{\rm N}$ hosts a zero energy reflectionless scattering mode (zero-RSM), namely $S_{\rm N}$ can be written as
$S_{\rm N}=\begin{pmatrix}r&t^{\prime}\\ t&r^{\prime}\end{pmatrix},$ (1)
where $r$ is a $m\times m$ matrix with $m<n$ and $\det(r^{0})=0$ with $r^{0}=r(E=0)$ [58]. We choose the SC phases in a way, such that the $m$ phases corresponding to the relevant states of the reflection matrix $r$ are zero, whereas the other $n-m$ phases are $\pi$. Then, it holds that
$e^{\imath\bm{\phi}}=\begin{pmatrix}I_{m}&0\\ 0&-I_{n-m}\end{pmatrix},$ (2)
with the $k\times k$ identitiy matrix $I_{k}$. Using Beenaker’s determinat formulas [60] at zero energy, $\det(\{S_{\rm N}^{0},e^{\imath\bm{\phi}}\})=0$ with $S_{\rm N}^{0}=S_{\rm N}(E=0)$, assuming time-reversal symmetry $S_{\rm N}=S_{\rm N}^{T}$, we find the simple expression
$\det\begin{pmatrix}r^{0}&0\\ 0&-{r^{\prime}}^{0}\end{pmatrix}=0,$ (3)
@@
which is fulfilled by assuming a zero-RSM. In addition, a reflectionless mode causes a transmission eigenvalue, i.e. an element of the spectrum $\sigma(tt^{\dagger})$, to be unity. Thus, in the presence of a zero-RSM, the MTJJ can be treated as an effective two terminal Josephson junction where a zero mode is reached by having a phase-shift of $\pi$ between the leads and unit transmission like is the case for a general two-terminal point contact [60].
So far, we have shown that zero-RSMs lead to zero energy solutions of ABSs. In the following, we will show that topological phase boundaries in multiterminal Josephson junctions arise from zero-RSMs.
@@
2-dot junction.– In the following we consider a double quantum dot junction that is tunnel coupled to four superconducting leads [59], as depicted in Fig. 1(a).
@@
Figure 1: (a) Four-terminal MTJJ with two dots with energies $\epsilon_{\rm L/R}$ coupled to four SC terminals with phases $\phi_{1,2,3,4}$ via couplings $\Gamma_{1,2,3,4}$ respectively. (b) Topological index $\Lambda(\epsilon,t)$ indicating which parts in parameter space may host non-trivial topology (black: $\Lambda=1$, white: $\Lambda=0$). Red solid lines (I) and (II) correspond to zero-RSMs of reflection matrix $r_{13}$ with $\epsilon\pm t=0$ ($E_{\pm}=0$). The dashed line (III) corresponds to another zero-RSM of the reflection matrix $r_{12}$ with $\epsilon^{2}-t^{2}+4=0$ ($E_{+}E_{-}+4=0$). (c) Self-energies $\Sigma_{1/2}$ in the complex plane for the topological phase boundaries (I) and (II) and for the non-trivial case (III). (d) The vectorfield $\bm{V}_{12}$ in Eq. 13 of the reflectionmatrix $r_{12}(z)$ with complex energy $z$ for the parameter configurations seen in the insets and in panel (b).
@@
Such models are used in the context of quantum dot physics [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], or as a tool to parametrize a scattering region with a few high-energy ABS levels [42, 31, 32, 53, 72, 51, 56]. We assume symmetry in the system, meaning that the on-site energies of the quantum dots and couplings are equal, $\epsilon\equiv\epsilon_{\rm L}=\epsilon_{\rm R}$ and $\Gamma\equiv\Gamma_{i}$ for $i=1,2,3,4$. We choose the phases such that $\phi_{1}=0$, $\bm{\phi}={\rm diag}(0,\phi_{2},\phi_{3},\phi_{4})$ and $\Gamma=1$. The low-energy physics of the system is described by an effective (topological) Hamiltonian [73, 74, 75]
@@
$\Sigma=\begin{pmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{pmatrix}=\Gamma\begin{pmatrix}1+e^{\imath\phi_{2}}&0\\ 0&e^{\imath\phi_{3}}+e^{\imath\phi_{4}}\end{pmatrix}$ (5)
@@
where $j\neq k$ and the total Chern number $C^{jk}=\sum_{i}C_{i}^{jk}$ as the sum over all occupied bands with the Berry curvature $B_{i}^{jk}\equiv-2{\rm Im}\braket{\partial_{\phi_{j}}v_{i}}{\partial_{\phi_{k}}v_{i}}$ of the $i$th eigenvector of $H_{\rm eff}$, |$\ket{v_{i}}$. We define the topological index $\Lambda$
@@
which is a measure for whether non-trivial topology can be achieved depending on the system parameters $\epsilon$ and $t$. In other words, if $\Lambda=1$, there is a finite non-trivial topological region in SC phase space $\{\bm{\phi}\}$.
In Fig. 1(b), we show the topological phase diagram $\Lambda(\epsilon,t)$ of the 2-dot junction which was calculated numerically using Ref. [76]. We observe three phase boundaries, indicated by the two straight lines (I) and (II) corresponding to the relation $\epsilon=\pm t$, and the curved line (III) with $\epsilon^{2}-t^{2}+4=0$. In particular, we find the following three aspects concerning the topological phase boundaries, i.e. they arise when:
@@
First, we consider the case where the self-energy in Eq. 5 is zero, $\Sigma=0$. This is the case when $\bm{\phi}=(0,\pi,\phi,\phi+\pi)$ with $\phi\in[0,2\pi)$. Then, it follows that $\det(H_{\rm eff})=\det(H_{\rm N})^{2}$ which is zero exactly when an eigenvalue of $H_{\rm N}$ is zero. However, how is that related to a reflectionless mode? Note, that the determinant of the reflection matrix $r$ can generally be written as a functional of $H_{\rm N}$ and the incoming $\Sigma_{\rm in}$ and outgoing $\Sigma_{\rm out}$ self-energies [58]. In particular, for the $2\times 2$ reflection matrix of lead 1 and 3, $r_{13}\equiv\begin{pmatrix}{S_{N}}_{1,1}&{S_{N}}_{1,3}\\ {S_{N}}_{3,1}&{S_{N}}_{3,3}\end{pmatrix}$, it holds that
@@
where $\Sigma_{\rm in}=\Sigma_{\rm out}=\Sigma_{0}/2$ and $\Sigma_{0}=\Sigma_{\bm{\phi}=0}$. It follows that $\det(H_{\rm eff})=0\Leftrightarrow\det(H_{\rm N})=0\Leftrightarrow\det(r_{13}^{0})=0$, exactly when $E_{\pm}=\epsilon\pm t=0$. Most interestingly, we find a relation between effective phase dependent self-energies $\Sigma$ [Eq. 5] and zero-RSMs of the reflection matrix $r_{13}$ which does not itself depend on the superconducting phases. To understand this connection, consider the following: No phase shift is assigned to incoming modes ($\Sigma_{\rm in}$) whereas a phase shift of $\pi$ is assigned to outgoing modes ($-\Sigma_{\rm out}$) as they leave the scattering region. Incoming and outgoing modes can be understood as gain and loss states, which in the language of non-hermitian theory is accounted for by the sign of the self-energy. This assignment of different phases is exactly analogous to the effect of SC couplings on the system, namely when the phases are varied to a phase difference of $\pi$. Thus, in symmetric junctions, zero-RSM effectively encode the same information as effective two-terminal Josephson junctions which are tuned to a $\pi$ phase difference.
Second, we consider the case where ${\rm tr}(\Sigma)=0$ with $\Sigma_{1}=-\Sigma_{2}\neq 0$ which results in the condition $\bm{\phi}=(0,\phi,\pi,\phi+\pi)$ for $\phi\in[0,2\pi)$. It holds that $-\Sigma_{2}=\Sigma_{1}=\gamma e^{\imath\alpha}$ for $\gamma\geq 0$ and $\alpha\in[0,2\pi)$. In the case of $\gamma=2$ [$\bm{\phi}=(0,0,\pi,\pi)$], it holds $\Sigma=2\sigma_{z}$ and
@@
which yields the condition $\epsilon^{2}-t^{2}+4=0$ or equivalently $E_{+}E_{-}+4=0$. Like before, we find the corresponding reflection matrix, which in this case is the $2\times 2$ matrix corresponding to terminal 1 and 2, $r_{12}\equiv\begin{pmatrix}{S_{N}}_{1,1}&{S_{N}}_{1,2}\\ {S_{N}}_{2,1}&{S_{N}}_{2,2}\end{pmatrix}$, where the incoming and outgoing self-energies read $\Sigma_{\rm in}=\sigma_{0}+\sigma_{z}$ and $\Sigma_{\rm out}=\sigma_{0}-\sigma_{z}$ resulting in
@@
It is then evident that $\det(H_{\rm eff})=0\Leftrightarrow\det(H_{\rm N})=0\Leftrightarrow\det(r_{12}^{0})=0$. As before, the reflection matrix imitates the effect of SC phases by assigning different phases to incoming and outgoing states respectively. In the previous case, incoming and outgoing states are mediated via both dots, meaning that $\Sigma_{\rm in}=\Sigma_{\rm out}$, which effectively eliminates any contribution from the self-energy in the numerator in Eq. 8. In this case, incoming and outgoing states see different self-energies, resulting in a non-trivial non-hermitian contribution in the numerator of Eq. 12. Hence, we get a more involved relation involving both energy levels granting a more complex phase boundary as seen in Fig. 1(b), (III). Furthermore, zero-RSMs are tied to geometric rules of the self-energies. This enables us to assign to each phase boundary observed in Fig. 1(b) a respective diagram like it is seen in Fig. 1(c).
Whereas the former considerations show that zero-RSMs provide ABSs at zero energy, it is not evident why they provide topological transitions. In order to demonstrate the relation to topology, consider a so-called reflection zero (R-zero) at a complex energy $z_{0}\in\mathbb{C}$, meaning that $\det(r(z_{0}))=0$. Such a R-zero provides a topological charge via a winding number in complex energy space [58]. To visualize this, we define the normalized vector field
@@
of the reflection matrix $r_{\alpha}(z)$ with complex energy $z\in\mathbb{C}$. In Fig. 1(d), we portray $\bm{V}_{12}$ of the reflection matrix $r_{12}(z)$ for the parameter configurations shown in the insets and also indicated in panel (b). For the left panel ($t=1.9$, $\epsilon=0$), it is seen that two R-zeros lie on opposite complex conjugated energies. The ”+”-symbols indicate topological charges resulting in a non-trivial winding number. In the middle panel ($t=2$, $\epsilon=0$), at the phase transition, two R-zeros combine to one zero-RSM at zero real energy. This point additionally corresponds to an exceptional point of the effective non-hermitian operator $z-H_{\rm N}-2\imath\sigma_{z}$. For the right panel ($t=2.1$, $\epsilon=0$) the single zero-RSM splits into two RSMs going away from zero real energy. We find that a topological phase boundary arises exactly when the topological singularity aligns with real zero energy.
3-dot junction.– In the following, we consider a three quantum dot junction depicted in Fig. 2(a).
Figure 2: (a) Four-terminal MTJJ with three dots with energies $\epsilon_{\rm L/M/R}$ coupled to four SC terminals with phases $\phi_{1,2,3,4}$ via couplings $\Gamma_{1,2,3,4}$ respectively. (b) Topological index $\Lambda(\epsilon,t)$ indicating which parts in parameter space may host non-trivial topology (black: $\Lambda=1$, white: $\Lambda=0$). Red solid lines (I) and (II) correspond to zero-RSMs of the reflection matrix $r_{145}$ with $\epsilon-t=0$ ($E_{-}=0$) and $\epsilon+2t=0$ ($E_{+}=0$). The dashed line (III) corresponds to the zero-RSM of $r_{123}$ with $\tilde{\epsilon}^{2}-\tilde{t}^{2}+4=0$. The dotted line (IV) corresponds to the topological phase boundary following $\hat{\epsilon}^{2}-\hat{t}^{2}+1=0$. (c) Self-energies $\Sigma_{1/2/3}$ in the complex plane for the zero-RSMs (I) and (II), for the non-trivial zero-RSM (III) and for the boundary not corresponding to a zero-RSM (IV). (d) The vectorfield $\bm{V}_{123}$ in Eq. 13 of the reflectionmatrix $r_{123}(z)$ for the parameter configurations seen in the insets and in panel (b).
Notably, this system was used in Ref. [56] to analyze spectroscopic data of a four terminal MTJJ setup. In addition, it was found in Ref. [56] that the model predicts non-trivial topology in a specific parameter regime. In the following, we show that we can attribute zero-RSMs of the normal scattering matrix to certain topological phase boundaries in the system. In addition, we find abnormal topological phase boundaries which have another, yet related, origin. The system is described by an effective Hamiltonian like in Eq. 4 with a normal Hamiltonian
$H_{\rm N}=\begin{pmatrix}\varepsilon&t&t\\ t&\varepsilon&t\\ t&t&\varepsilon\end{pmatrix}$ (14)
and the self-energy matrix
$\Sigma=\begin{pmatrix}\Sigma_{1}&0&0\\ 0&\Sigma_{2}&0\\ 0&0&\Sigma_{3}\end{pmatrix}=\begin{pmatrix}1+e^{\imath\phi_{2}}&0&0\\ 0&e^{\imath\phi_{2}}+e^{\imath\phi_{3}}&0\\ 0&0&e^{\imath\phi_{3}}+e^{\imath\phi_{4}}\end{pmatrix}.$ (15)

### v3

@@ Abstract
Multiterminal Josephson junctions are a promising platform to study non-trivial topology in engineered quantum systems. Yet, experimentally meaningful insight into what exactly makes these systems topologically non-trivial remains elusive. In this work, we show that zero energy reflectionless scattering modes (RSMs) of the normal scattering matrix result in topological phase boundaries. By analyzing two different setups, we explain the origin of each topological phase boundary and provide generalizable insight into these systems. The considerations here can be of help for experimentalists as it connects the properties of the normal scattering region to the Andreev bound state spectrum of multiterminal superconducting junctions. In addition, our work provides a way to give meaning to the edge modes of these engineered topological systems with synthetic dimensions by providing a correspondence between unity transmission modes and boundaries between topological trivial and non-trivial phases.
@@
Introduction.– Topology manifests itself as one of the most fundamental aspects of condensed matter systems. Such systems include topological insulators [Kane2005, Kane2005a, Bernevig2006, Fu2007, Konig2007], topological superconductors [Kitaev2001, Fu2008, Sato2009, Alicea2012] and quantum Hall systems [klitzing1980new, TKKN, hatsugai1993chern]. Another promising direction lies in the ability to engineer the topological properties via control parameters to study topological phase transitions. Examples include topological photonics [lu2014topological, wang2009observation, khanikaev2013photonic], topological-driven Floquet systems [rudner2020band, kitagawa2010topological], topological electrical [imhof2018topolectrical], and superconducting circuits [fatemi2021weyl, peyruchat2021transconductance, herrig2022cooper, Peyruchat2024a]. Furthermore, while the existence of a bulk-boundary correspondence in such topological synthetic systems is often not evident; it is still possible to infer effective boundary effects from bulk topology with time-boundary effects [Xu2025] or engineered boundaries [Dutt2022]. Multiterminal Josephson junctions (MTJJs) have proven to be another promising route towards engineered topology where the topology is encoded in the superconducting phase degree of freedom [riwar2016multi, eriksson2017topological, xie2017topological, xie2018weyl, deb2018josephson, xie2019topological, gavensky2019topological, houzet2019majorana, klees2020microwave, klees2021ground, Weisbrichtensor, PhysRevB.107.165301, Peralta_Gavensky_2023, riwar2019fractional, PhysRevB.107.035408, weisbrich2021second, xie2022non, Yokoyama2017, Yokoyama2015a, Meyer2017b, Repin2022, Barakov2023a]. Additionally, MTJJs also host novel (topological) phases in the context of non-hermiticity [cayao2023bulk, PhysRevLett.133.086301, cayao2023nonhermitian, PhysRevB.109.214514, 10.1063/5.0215522, Pino2024, cayao2024nonhermitianminimalkitaevchains, cayao2024nonhermitianmultiterminalphasebiasedjosephson, Ohnmacht2024a]. Whereas observable consequences of the non-trivial topology are predicted to exist [riwar2016multi, klees2020microwave], whether these signatures can be observed in real systems is not clear [Repin2019a]. Experimentally, recent advances in sample fabrication lead to the spectroscopic measurement of Andreev bound states (ABSs) with phase control via flux lines [Coraiola2023b] which also represents an experimental realization of an Andreev molecule [Pillet2019a, Kornich2019a]. In particular, phase manipulation of a four-terminal MTJJ was established experimentally in Ref. [Antonelli2025], which satisfies the minimal requirement for the existence of Weyl nodes in MTJJs with time-reversal symmetry [riwar2016multi]. Whereas the theory provided in Ref. [Antonelli2025] predicts non-trivial topology in accordance with spectroscopic data, additional experimental evidence is needed to proof the existence of non-trivial topology.
@@
Although MTJJs and their topology have been studied in detail for nearly a decade, making a connection from non-trivial topology to experimentally relevant parameters is difficult. Promising samples must host ABSs very close to zero energy, yet in a multiterminal setup, such a zero energy ABS is not necessarily related to a unity transmission mode as it is not possible to label ABSs by transmission eigenvalues of the full scattering matrix [VanHeck2014]. Furthermore, a bulk-boundary correspondence in these systems still remains elusive, as there is no clear analog of a boundary in topological systems with synthetic dimensions.
@@
In this manuscript, we provide a concrete explanation for the appearance of non-trivial topology and Weyl nodes in MTJJs by demonstrating that reflectionless scattering modes (RSMs) [PhysRevA.102.063511] of the normal scattering matrix result in topological phase boundaries. We illustrate our findings by analyzing a four-terminal hybrid system used in Ref. [10.21468/SciPostPhys.15.5.214] where we discuss the conditions for topological phase boundaries in detail. We show that each phase boundary is related to a zero-energy RSM (zero-RSM) of the normal scattering matrix. In the second part, we analyze the four-terminal system used in Ref. [Antonelli2025] and its corresponding topological phase diagram in detail. We find that most topological phase boundaries correspond to zero-RSMs of the normal scattering matrix. Additionally, we find abnormal topological phase boundaries, which do not arise from RSMs, yet are related to effective RSMs. Additionally, we find that the source of topology is inherent to the non-trivial topology of RSMs [PhysRevA.102.063511]. The findings here help to provide explanations for the occurrence of zero-energy solutions and topology in experimental MTJJ setups. Furthermore, the topological phase boundaries defined in this work provide a way to define a boundary of the synthetic bulk, thus defining an effective bulk-boundary correspondence with unity transmission modes.
@@
Reflectionless scattering modes (RSMs).– We consider a normal scattering region parametrized by a $n\times n$ normal scattering matrix $S_{\rm N}$, coupled to $n$ superconducting (SC) leads with phases $\bm{\phi}={\rm diag}(\phi_{1},...,\phi_{n})$. Let’s assume that $S_{\rm N}$ hosts a zero energy reflectionless scattering mode (zero-RSM). Then, $S_{\rm N}=\begin{pmatrix}r&t^{\prime}\\ t&r^{\prime}\end{pmatrix}$ with the $m\times m$ ($m<n$) reflection matrix $r$ such that $\det(r^{0})=0$ with $r^{0}=r(E=0)$ [PhysRevA.102.063511]. We choose the SC phases such that the $m$ phases corresponding to the relevant states of the reflection matrix $r$ are zero, whereas the other nmn-m phases are $\pi$. Then, it holds that $e^{\imath\bm{\phi}}={\rm diag}(I_{m},-I_{n-m})$ with the $k\times k$ identity matrix $I_{k}$. Using Beenakker’s determinant formulas [PhysRevLett.67.3836] at zero energy, $\det(\{S_{\rm N}^{0},e^{\imath\bm{\phi}}\})=0$ with $S_{\rm N}^{0}=S_{\rm N}(E=0)$, assuming time-reversal symmetry $S_{\rm N}=S_{\rm N}^{T}$, we find the simple expression $\det{\rm diag}(r^{0},-{r^{\prime}}^{0})=0$ which is fulfilled by assuming a zero-RSM. In addition, a RSM causes a transmission eigenvalue, i.e. an element of the spectrum $\sigma(tt^{\dagger})$, to be unity. Thus, in the presence of a zero-RSM, the MTJJ can be treated as an effective two terminal Josephson junction where a zero mode is reached by having a phase-shift of $\pi$ between the leads and unit transmission like is the case for a general two-terminal point contact [PhysRevLett.67.3836].
@@
2-dot junction.– We consider a double quantum dot junction that is tunnel coupled to four superconducting leads [10.21468/SciPostPhys.15.5.214], as depicted in Fig. 1(a).
@@
Figure 1: (a) Four-terminal MTJJ with two dots with energies $\epsilon_{\rm L/R}$ coupled to four SC terminals with phases $\phi_{1,2,3,4}$ via couplings $\Gamma_{1,2,3,4}$. (b) Topological index $\Lambda(\epsilon,t)$ indicating which parts in parameter space may host non-trivial topology. Red solid lines (I) and (II) correspond to zero-RSMs of reflection matrix $r_{13}$ with $\epsilon\pm t=0$ ($E_{\pm}=0$). The dashed line (III) corresponds to another zero-RSM of the reflection matrix $r_{12}$ with $\epsilon^{2}-t^{2}+4=0$ ($E_{+}E_{-}+4=0$). (c) Self-energies $\Sigma_{1/2}$ in the complex plane for the topological phase boundaries (I) and (II) and for the non-trivial case (III). (d) Four-terminal MTJJ with three dots with energies $\epsilon_{\rm L/M/R}$ coupled to four SC terminals with phases $\phi_{1,2,3,4}$. (e) Topological index $\Lambda(\epsilon,t)$ as in panel (b). Red solid lines (I) and (II) correspond to zero-RSMs of the reflection matrix $r_{145}$ with $\epsilon-t=0$ ($E_{-}=0$) and $\epsilon+2t=0$ ($E_{+}=0$). The dashed line (III) corresponds to the zero-RSM of $r_{123}$ with $\tilde{\epsilon}^{2}-\tilde{t}^{2}+4=0$. The dotted line (IV) corresponds to the topological phase boundary following $\hat{\epsilon}^{2}-\hat{t}^{2}+1=0$. (f) Self-energies $\Sigma_{1/2/3}$ in the complex plane for the zero-RSMs (I) and (II), for the non-trivial zero-RSM (III) and for the boundary not corresponding to a zero-RSM (IV). (g) The vectorfield $\bm{V}_{12}$ in Eq. 10 of the reflection matrix $r_{12}(z)$ with complex energy $z$ for the parameter configurations seen in the insets and in panel (b). (h) The vectorfield $\bm{V}_{123}$ of the reflection matrix $r_{123}(z)$ for the parameter configurations seen inthe insets in panel (e).
@@
Such models are used in the context of quantum dot physics [martin2011josephson, recher2001andreev, governale2008real, pala2007nonequilibrium, eldridge2010superconducting, futterer2009nonlocal, herrmann2010carbon, tanaka2010correlated, tanaka2008andreev, jonckheere2009nonequilibrium, de2010hybrid], or as a tool to parametrize a scattering region with a few high-energy ABS levels [Meyer2017b, klees2020microwave, klees2021ground, Coraiola2023b, Ohnmacht2024, Ohnmacht2024a, Antonelli2025]. We assume symmetry in the system, meaning that the on-site energies of the quantum dots and couplings are equal, $\epsilon\equiv\epsilon_{\rm L/R}$ and $\Gamma\equiv\Gamma_{i}$ for $i=1,2,3,4$. We choose the phases such that $\phi_{1}=0$, $\bm{\phi}={\rm diag}(0,\phi_{2},\phi_{3},\phi_{4})$ and $\Gamma=1$. The low-energy physics of the system is described by an effective (topological) Hamiltonian [Wang2012, Wang2013, Gavensky2023]
@@
$\Sigma=\begin{pmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{pmatrix}=\begin{pmatrix}1+e^{\imath\phi_{2}}&0\\ 0&e^{\imath\phi_{3}}+e^{\imath\phi_{4}}\end{pmatrix}$ (2)
@@
which is a measure for whether non-trivial topology can be achieved depending on the system parameters $\epsilon$ and $t$. In other words, if $\Lambda=1$, there is a finite non-trivial topological region in SC phase space $\{\bm{\phi}\}$. Furthermore, we notice that a boundary between a topologically non-trivial ($\Lambda\neq 0$) and trivial region ($\Lambda=0$) can be viewed as a boundary of a bulk topological phase.
In Fig. 1(b), we show the topological phase diagram $\Lambda(\epsilon,t)$ of the 2-dot junction which was calculated numerically using Ref. [Fukui2005]. We observe three phase boundaries, indicated by the two straight lines (I) and (II) corresponding to the relation $\epsilon=\pm t$, and the curved line (III) with $\epsilon^{2}-t^{2}+4=0$. In particular, we find the following three aspects concerning the topological phase boundaries, i.e. they arise when:
@@
First, we consider the case where the self-energy in Eq. 2 is zero, $\Sigma=0$. This is the case when $\bm{\phi}=(0,\pi,\phi,\phi+\pi)$ with $\phi\in[0,2\pi)$. Then, it follows that $\det(H_{\rm eff})=\det(H_{\rm N})^{2}$ which is zero exactly when an eigenvalue of $H_{\rm N}$ is zero. However, how is that related to a RSM? Note, that the determinant of the reflection matrix $r$ can generally be written as a functional of $H_{\rm N}$ and the incoming $\Sigma_{\rm in}$ and outgoing $\Sigma_{\rm out}$ self-energies [PhysRevA.102.063511]. In particular, for the $2\times 2$ reflection matrix of lead 1 and 3, $(r_{13})_{ij}=(S_{\rm N})_{ij}$ for $i,j=1,3$, it holds that
@@
where $\Sigma_{\rm in}=\Sigma_{\rm out}=\Sigma_{0}/2$ and $\Sigma_{0}=\Sigma_{\bm{\phi}=0}$. It follows that $\det(H_{\rm eff})=0\Leftrightarrow\det(H_{\rm N})=0\Leftrightarrow\det(r_{13}^{0})=0$, exactly when $E_{\pm}=\epsilon\pm t=0$. Most interestingly, we find a relation between effective phase dependent self-energies $\Sigma$ [Eq. 2] and zero-RSMs of the reflection matrix $r_{13}$ which does not itself depend on the superconducting phases. To understand this connection, consider the following in the normal state: No phase shift is assigned to incoming modes ($\Sigma_{\rm in}$) whereas a phase shift of $\pi$ is assigned to outgoing modes ($\Sigma_{\rm out}$) as they leave the scattering region. Incoming and outgoing modes can be understood as gain and loss states, which in the language of non-hermitian theory is accounted for by the sign of the self-energy. This assignment of different phases is exactly analogous to the effect of SC couplings on the system, namely when the phases are varied to a phase difference of $\pi$. Thus, in symmetric junctions, zero-RSMs effectively encode the same information as effective two-terminal Josephson junctions which are tuned to a $\pi$ phase difference.
Now, let’s consider the case where ${\rm tr}(\Sigma)=0$ with $\Sigma_{1}=-\Sigma_{2}\neq 0$ which results in the condition $\bm{\phi}=(0,\phi,\pi,\phi+\pi)$ for $\phi\in[0,2\pi)$. It holds that $\Sigma_{2}=\Sigma_{1}=\gamma e^{\imath\alpha}$ for $\gamma\geq 0$ and $\alpha\in[0,2\pi)$. In the case of $\gamma=2$ $\bm{\phi}=(0,0,\pi,\pi)$], it holds $\Sigma=2\sigma_{z}$ and
@@
which yields the condition $\epsilon^{2}-t^{2}+4=0$ or equivalently $E_{+}E_{-}+4=0$. Like before, we find the corresponding reflection matrix, which in this case is the $2\times 2$ matrix corresponding to terminal 1 and 2, $(r_{12})_{ij}=(S_{\rm N})_{ij}$ for $i,j=1,2$, where the incoming and outgoing self-energies read $\Sigma_{\rm in}=\sigma_{0}+\sigma_{z}$ and $\Sigma_{\rm out}=\sigma_{0}-\sigma_{z}$ resulting in
@@
It is then evident that $\det(H_{\rm eff})=0\Leftrightarrow\det(r_{12}^{0})=0$. As before, the reflection matrix imitates the effect of SC phases by assigning different phases to incoming and outgoing states respectively. In the previous case, incoming and outgoing states are mediated via both dots, meaning that $\Sigma_{\rm in}=\Sigma_{\rm out}$, which effectively eliminates any contribution from the self-energy in the numerator in Eq. 5. In this case, incoming and outgoing states see different self-energies, resulting in a non-trivial non-hermitian contribution in the numerator of Eq. 9. Hence, we get a more involved relation involving both energy levels granting a more complex phase boundary as seen in Fig. 1(b), (III). Furthermore, zero-RSMs are tied to geometric rules of the self-energies. This enables us to assign to each phase boundary observed in Fig. 1(b) a respective diagram like it is seen in Fig. 1(c).
Whereas the former considerations show that zero-RSMs provide ABSs at zero energy, it is not evident why they provide topological transitions. In order to demonstrate the relation to topology, consider a so-called reflection zero (R-zero) at a complex energy $z_{0}\in\mathbb{C}$, meaning that $\det(r(z_{0}))=0$. Such a R-zero provides a topological charge via a winding number in complex energy space [PhysRevA.102.063511]. To visualize this, we define the normalized vector field
@@
of the reflection matrix $r_{\alpha}(z)$ with complex energy $z\in\mathbb{C}$. In Fig. 1(g), we portray $\bm{V}_{12}$ of the reflection matrix $r_{12}(z)$ for the parameter configurations shown in the insets and also indicated in panel (b). For the left panel ($t=1.9$, $\epsilon=0$), it is seen that two R-zeros lie on opposite complex conjugated energies. The ”+”-symbols indicate topological charges resulting in a non-trivial winding number. In the middle panel ($t=2$, $\epsilon=0$), at the phase transition, two R-zeros combine to one zero-RSM at zero real energy, forming an exceptional point of the effective non-hermitian operator $z-H_{\rm N}-2\imath\sigma_{z}$. For the right panel ($t=2.1$, $\epsilon=0$) the single zero-RSM splits into two RSMs going away from zero real energy. We find that a topological phase boundary arises exactly when the topological singularity of R-zeros aligns with real zero energy.
3-dot junction.– In the following, we consider a three quantum dot junction depicted in Fig. 1(d). Notably, this system was used in Ref. [Antonelli2025] to analyze spectroscopic data of a four-terminal MTJJ. In addition, it was found in Ref. [Antonelli2025] that the model predicts non-trivial topology in a specific parameter regime. In the following, we show that we can attribute zero-RSMs of the normal scattering matrix to certain topological phase boundaries in the system. In addition, we find abnormal topological phase boundaries which have another, yet related, origin. The system is described by an effective Hamiltonian like in Eq. 1 with a normal Hamiltonian and the self-energy matrix
$H_{\rm N}$ $=\begin{pmatrix}\varepsilon&t&t\\ t&\varepsilon&t\\ t&t&\varepsilon\end{pmatrix}$, ($1$)
$\Sigma$ $={\rm diag}(1+e^{\imath\phi_{2}},e^{\imath\phi_{2}}+e^{\imath\phi_{3}},e^{\imath\phi_{3}}+e^{\imath\phi_{4}}).$ (12)